# PO-Flow: Flow-based Generative Models for Sampling Potential Outcomes and Counterfactuals

### v2

@@ Abstract
Predicting potential and counterfactual outcomes from observational data is central to clinical decision-making, where physicians must weigh treatments for an individual patient rather than relying solely on average effects at the population level. We propose PO-Flow, a continuous normalizing flow (CNF) framework for causal inference that jointly models potential outcomes and counterfactuals. Trained via flow matching, PO-Flow provides a unified approach to average treatment effect estimation, individualized potential outcome prediction, and counterfactual prediction. Besides, PO-Flow directly learns the densities of potential outcomes, enabling likelihood-based evaluation of predictions. Furthermore, PO-Flow explores counterfactual outcome generation conditioned on the observed factual in general observational datasets, with a supporting recovery result under certain assumptions. PO-Flow outperforms modern baselines across diverse datasets and causal tasks in the potential outcomes framework.
1 Introduction
Predicting potential outcomes (POs) lies at the core of causal inference. In a typical setting, each individual has two potential outcomes: $Y^{(1)}$ if they receive a treatment ($A=1$) and $Y^{(0)}$ if they do not ($A=0$). However, in the real world, only one outcome is observed, while the other potential outcome remains fundamentally unknown [25]. Accurately estimating both POs is crucial for decision-making in healthcare [37, 24, 46], where clinicians may wish to compare a patient’s likely outcomes under different treatment options rather than rely solely on average effects [56, 17].
In the Potential Outcomes (POs) literature, the vast majority of methods focus on estimating the conditional average treatment effect (CATE): $\mathbb{E}[Y^{(1)}-Y^{(0)}|X]$, at the population level [28, 22, 1, 19, 33, 38, 50, 61]. It is important to note that although some methods perform well in estimating population-level CATE, they may underperform in individualized POs prediction, as accurately predicting POs in finite samples is challenging [47].
@@ 1 Introduction
However, making decisions for individuals based on average treatment effects can be misleading. Therefore, accurately predicting individualized POs and learning their full distributions $p(Y^{(A)}|X,A)$ is a crucial task [63, 31]. Recent works [61, 40, 41] have explored generative models to sample POs and estimate empirical distributions. However, methods that learn the full density of POs without relying on explicit distributional assumptions (e.g., Gaussian mixtures) remain scarce.
In addition, another critical challenge is estimating the counterfactual outcome conditioned on the factual: $Y_{\mathrm{CF}}^{(1-A)}:=Y^{(1-A)}|X,Y^{(A)}$, i.e., predicting what would have happened under the alternative treatment, given the observed outcome $Y^{(A)}$. This task is especially relevant in post-hoc clinical analyses, where practitioners ask, “Would the patient have fared better on another therapy?” [60, 18].
Counterfactual prediction differs from standard PO estimation—it requires reasoning about outcomes under an alternative treatment conditioned on the observed factual outcome, introducing dependencies between observed and unobserved outcomes. Structural Causal Models (SCMs) [44] and modern extensions [42, 6, 12, 7, 59, 45, 8] offer the most widely used theoretical framework for counterfactual reasoning. SCMs methods typically assume known causal structures or aim to learn them explicitly. However, it may be possible to estimate counterfactuals directly without recovering the SCM [64]. In PO-Flow, we do not explicitly assume an SCM, but seek to directly estimate counterfactual outcomes.
@@
Despite recent progress, two key gaps remain in the potential outcomes literature. First, counterfactual outcome estimation conditioned on the observed factual outcome has received little attention, though it is crucial for post-hoc analyses. Second, beyond counterfactuals, most methods are specialized—some target CATE estimation, others aim at potential outcome prediction or distributional modeling—but few provide a unified and scalable algorithmic framework that can jointly address these tasks.
1.1 Main Contributions
We propose PO-Flow, a continuous normalizing flow (CNF) framework for potential outcomes and counterfactuals. Our main contributions are: (1) Enable counterfactual outcome generation conditioned on the observed factual, together with a supporting counterfactual recovery result (Corollary 4.5) under certain assumptions; (2) Provide a unified generative framework for potential outcome prediction, CATE estimation, and general-purpose counterfactual prediction; (3) Learn the density of potential outcomes, supporting principled uncertainty quantification; (4) Achieve superior performance over state-of-the-art methods across diverse causal tasks and benchmark datasets.
1.2 Related Works
CATE Estimation: Numerous methods have been proposed for CATE estimation [28, 22, 1, 19, 33, 38, 50, 61]. Broadly, these fall into model-agnostic and model-specific approaches. Model-agnostic estimators, or meta-learners [33, 15, 14, 2, 49], typically follow two strategies [39]: (a) one-step methods, which fit regression models and compute CATE as the difference in predicted outcomes, and (b) two-step methods, which estimate nuisance functions to construct pseudo-outcomes, then regress them on covariates.
In contrast, model-specific estimators adapt machine learning techniques for CATE estimation, including methods such as [13, 50, 38, 30, 61]. More recently, neural networks have been widely applied to this task, with notable examples in [50, 38, 61]. However, most of these approaches mainly focus on estimating CATE rather than predicting individualized potential outcomes.
Individualized POs Prediction: Although some CATE methods can be adapted for potential outcomes (POs) prediction, they may perform suboptimally, particularly on finite samples, as they are not specifically optimized for accurate PO estimation. This limitation partly arises because many methods assume that CATE has a simpler functional structure than POs. Nevertheless, several CATE approaches are also designed with POs prediction as the target [33, 50, 61, 38, 41, 39].
Counterfactual Outcomes Prediction: Structural Causal Models (SCMs) provide a well-established framework for counterfactual reasoning [44, 42, 6, 12, 7, 59, 45, 8], but practical applications often require specifying the full causal graph and structural equations, which may be difficult or infeasible for real-world data. Some recent works avoid explicitly modeling SCMs [35, 5, 40, 57], but typically address specific tasks and rely on additional information, such as temporal structure.
Some works [24, 30, 38, 34, 61] targeting CATE estimation describe their goal as counterfactual inference. However, in these works, "counterfactual" typically refers to the unobserved potential outcome under an alternative treatment, without conditioning on the observed factual outcome.
While SCMs offer theory-grounded counterfactual reasoning, general-purpose methods for individualized counterfactual prediction—without explicitly modeling SCMs or relying on auxiliary information—remain scarce. Although [61, 39] were not originally designed for counterfactuals, we can adapt them by incorporating factual information into intermediate steps and use them as baselines.
Diffusion and Flow Models: Diffusion models [23, 52, 54] and flow-based generative models [10, 58, 36, 4] have demonstrated strong performance across a range of applications. Recently, diffusion and flow-based generative models have been applied to causal inference tasks such as counterfactual image generation [48, 16], causal query answering [9], and inference over causal graphs or structural causal models [8, 29]. To our knowledge, two generative algorithms have been proposed specifically for the potential outcomes (PO) task: the diffusion-based DiffPO [39] and the discrete normalizing flow-based INFs [41]. We conduct extensive comparisons with both methods in this work.
2 Preliminaries
2.1 Neural ODE and Continuous Normalizing Flow:
A Neural ODE is a continuous model where the trajectory of data is modeled as the solution of an ordinary differential equation (ODE). Formally, in $\mathbb{R}^{d}$, given an input $y_{0}=y(0)$ at time $t=0$, the transformation to the output $y_{1}=y(1)$ at time $t=1$ is governed by:
$\small\frac{dy(t)}{dt}=v_{\theta}(y(t),t),\quad t\in[0,1],$ (1)
@@ 2.1 Neural ODE and Continuous Normalizing Flow:
where $v_{\theta}:\mathbb{R}^{d}\times[0,1]\to\mathbb{R}^{d}$ represents the velocity field, which is parameterized by a neural network with parameter $\theta$. The time horizon is rescaled to $[0,1]$ without loss of generality.
Continuous Normalizing Flow (CNF) is a class of normalizing flows in which the transformation of a probability density is governed by a time-continuous Neural ODE. The marginal density of $y(t)$, denoted as $p(y,t)$, evolves according to the continuity equation implied by the ODE in Eq. (1). This continuity equation is given by:
$\small\partial_{t}p(y,t)+\nabla\cdot(p(y,t)v_{\theta}(y,t))=0,\quad t\in[0,1],$ (2)
where $\nabla\cdot$ denotes the divergence operator.
When the Neural ODE is well-posed, it defines a continuous and invertible mapping from the initial sample $y_{0}$ to the terminal sample $y_{1}$. The inverse mapping from $y_{1}$ to $y_{0}$ can be computed by integrating the ODE (1) in reverse time.
This framework allows setting $y_{0}\sim p(\cdot,0)$ as the data distribution and $y_{1}\sim p(\cdot,1)$ as the base (noise) distribution, typically chosen as $N(0,I)$. For notational convenience, we define $p_{Y}(\cdot):=p(\cdot,0)$ as the data distribution and $q(\cdot):=p(\cdot,1)$ as the base (noise) distribution. Throughout this paper, we adopt the common choice $q(\cdot)=N(0,I)$.
2.2 Flow Matching
Flow Matching (FM) [36, 4] provides a simulation-free approach for training continuous normalizing flows by directly regressing the model velocity $v_{\theta}(y,t)$ onto a predefined target velocity field $u(y,t)$. The original FM objective is defined as:
@@ 2.2 Flow Matching
$\small\mathcal{L}_{\text{FM}}(\theta)=\mathbb{E}_{t\sim\mathcal{U}[0,1],y\sim p(\cdot,t)}\left[\|v_{\theta}(y,t)-u(y,t)\|^{2}\right],$ (3)
where $u(y,t)$ is an analytically specified target velocity field.
Linear Interpolant: Due to the intractability of $u(y,t)$ without conditioning on the initial data $y_{0}$, Conditional Flow Matching (CFM) is introduced. Conditioned on an observed sample $y_{0}\sim p_{Y}(\cdot)$, CFM employs a pre-specified interpolation function $I_{t}$ between two endpoints $y_{0}$ and $y_{1}\sim q(\cdot)$, where $q$ is typically $N(0,I)$. The interpolated path is defined as:
$\small\phi(y_{0},y_{1};t):=I_{t}(y_{0},y_{1}),\quad t\in[0,1].$ (4)
Here, $I_{t}$ can be analytically specified. Conditioned on $y_{0}$, the optimal and most information-preserving choice is the linear interpolation path [36]:
$\small\phi(y_{0},y_{1};t)=(1-t)y_{0}+(t+\sigma_{\text{min}}(1-t))y_{1},$ (5)
where $y_{1}\sim N(0,I)$ and $\sigma_{\text{min}}$ is a small positive hyperparameter that ensures $p(\phi,0)\sim N(y_{0},\sigma_{\text{min}}^{2}I)$. When $\sigma_{\text{min}}=0$, this reduces to the strict linear interpolant $\phi=(1-t)y_{0}+ty_{1}$.
Here, $\phi$ serves as a pre-specified reference trajectory that the ODE solution $y(t)$ should follow at time $t$. Under the linear interpolant defined in (5), the corresponding reference velocity field becomes:
$\small\frac{d\phi}{dt}=(1-\sigma_{\text{min}})y_{1}-y_{0}.$ (6)
Training loss: CFM trains the flow $v_{\theta}(y(t),t)$ by directly regressing it onto the reference velocity field (6). The training objective is given by:
$\small\mathcal{L}_{\text{CFM}}(\theta)=\ \mathbb{E}_{t\sim\mathcal{U}[0,1],y_{0}\sim p_{Y}(\cdot),y_{1}\sim q(\cdot)}\|v_{\theta}(\phi,t)-\frac{d\phi}{dt}\|^{2}.$ (7)
3 Problem Settings
Notations and Setups: We consider a setting where each individual is described by covariates $X\in\mathbb{R}^{d_{X}}$, a treatment assignment $A\in\{0,1\}$, and a potential outcome $Y^{(A)}\in\mathbb{R}$ under treatment $A$. The observational dataset is $\mathcal{D}=\{(y^{(a)}_{i},x_{i},a_{i})\}_{i=1}^{n}\sim p_{Y,X,A}(\cdot)$, where $y_{i}^{(a)}$ is the observed potential outcome under treatment $a_{i}$, and $(x_{i},a_{i})\sim p_{X,A}(\cdot)$ are the associated covariates and treatment.
@@ 3 Problem Settings
Potential Outcomes: In the potential outcomes framework, each individual has two potential outcomes: $Y^{(0)}$ (control) and $Y^{(1)}$ (treatment), but only one is observed based on assignment $A$. We define potential outcome prediction as estimating $Y^{(a)}\mid X,A=a$. For convenience, we denote $Y^{(a)}$ and $\hat{Y}^{(a)}$ as the true and predicted potential outcomes.
Counterfactual Outcomes: In contrast to potential outcomes, we define the counterfactual outcome as $Y_{\text{cf}}^{(1-a)}:=Y^{(1-a)}\mid X,A=a,Y^{(a)}$, which is the outcome had the individual received the opposite treatment, conditioned on its observed potential (factual) outcome. For convenience, we denote $Y_{\text{cf}}^{(1-a)}$ and $\hat{Y}_{\text{cf}}^{(1-a)}$ as the true and predicted counterfactual outcomes, respectively.
@@
To ensure the identifiability of potential outcomes (in Section 4.1), we adopt the standard assumptions commonly used in the literature [39, 41, 61]:
Assumption 3.1.
(1) Consistency: If an individual receives treatment $A=a$, the observed outcome equals the corresponding potential outcome: $Y=Y^{(a)}$. (2) Unconfoundedness: All confounders are observed, i.e., $\{Y^{(0)},Y^{(1)}\}\perp\!\!\!\perp A\mid X$. (3) Overlap: Treatment assignment is probabilistic: $0<P(A=1\mid X=x)<1$, for all $x\in\mathcal{X}$.
@@
In Section 4.2, we will introduce additional assumptions on counterfactual outcomes estimation in PO-Flow and provide the corresponding theoretical analyses for counterfactual recovery.
4 PO-Flow: A Continuous Normalizing Flow for Potential Outcomes
We introduce PO-Flow, a continuous normalizing flow framework for modeling potential and counterfactual outcomes, with three main goals: 1. predicting individualized potential outcomes, 2. predicting counterfactual outcomes, and 3. learning the densities.
Our PO-Flow framework is built on Continuous Normalizing Flows (CNFs) and is trained using the Flow Matching (FM) objective. Our learning target is the trajectory of the sample $y(t)$ over $t\in[0,1]$. To condition the dynamics on both covariates $x$ and treatment $a$, the Neural ODE is defined as:
$\small\frac{dy(t)}{dt}=v_{\theta}(y(t),t;x,a),\quad t\in[0,1],$ (8)
where $v_{\theta}$ is parametrized by a neural network, with $y(0)$ following the data distribution and $y(1)\sim q(\cdot)$.
Notation: Given a data point $(y^{(a)},x,a)\sim p_{Y,X,A}(\cdot)$, we define $y_{0}=y^{(a)}$ as the observed training sample at time $t=0$ in the Neural ODE, and $y_{1}\sim q(\cdot)$ as the noise sample at time $t=1$. Recall that we denote $\hat{y}^{(a)}$ and $\hat{y}_{\text{cf}}^{(1-a)}$ as the predicted potential and counterfactual outcomes, respectively.
Training loss: PO-Flow adopts the Conditional Flow Matching (CFM) objective for training. Under the updated definitions, our training loss retains the form of Equation (7) and is given by:
$\min_{\theta}\ \mathcal{L}_{\text{CFM}}(\theta)=\ \mathbb{E}_{t\sim\mathcal{U}[0,1],p_{Y,X,A}(\cdot),q(\cdot)}\|v_{\theta}(\phi,t;x,a)-\frac{d\phi}{dt}\|^{2},$ (9)
where $\phi$ and $\frac{d\phi}{dt}$ are defined in (5) and (6), respectively, with $y_{0}=y^{(a)}$ and $y_{1}\sim q(\cdot)$.
Remark: Following prior works [20, 39], one may estimate propensity scores $w(x)=p(A=1\mid X=x)$ with a pre-trained classifier and apply inverse weighting $\frac{1}{w(x)}$ in the loss. We found this provided little performance improvement, but both implementations are included in our code.
Forward Process: We treat the forward process as an encoding operation, denoted by the function $Z:=\Phi_{\theta}(Y^{(A)};\,X,A)$. Given the velocity field $v_{\theta}$, the forward process refers to pushing a factual outcome $y^{(a)}$ (at $t=0$) to a latent embedding $z$ (at $t=1$) via:
$z=\Phi_{\theta}(y^{(a)};\,x,a)=y^{(a)}+\int_{0}^{1}v_{\theta}(y(t),t;x,a)dt,\ \text{with}\ y(0)=y^{(a)}.$ (10)
Through the forward process, $z$ encodes information from the factual outcome $y^{(a)}$ as well as its associated covariates $x$ and treatment $a$.
Reverse Process: We treat the reverse process as a decoding operation. It is initialized differently for potential outcome (Section 4.1) and counterfactual prediction (Section 4.2). In general, at $t=1$, given a latent representation $z$, which can be either sampled from $q(\cdot)=N(0,1)$ or obtained by encoding a factual sample, the reverse process is defined as:
$\hat{y}_{0}=\Phi_{\theta}^{-1}(z;x,a)=z-\int_{0}^{1}v_{\theta}(y(t),t;x,a)dt,\ \text{with}\ y(1)=z.$ (11)
Empirically, we use the Runge-Kutta method (Appendix B.3) for numerical integration.
During training, we construct synthetic paths $\phi$ between $y_{0}$ and $y_{1}\sim q(\cdot)$ using the CFM loss conditioned on $y_{0}$. As shown in Proposition A.1 (Appendix), the gradient of the CFM loss is equivalent to that of the original FM loss. Consequently, the learned velocity field $v_{\theta}$ generates $\hat{y}_{0}$ independent of the specific training samples $y_{0}$, and since $y_{1}\sim q(\cdot)$ is random, $v_{\theta}$ induces a distribution over outcomes conditioned on $(x,a)$.
4.1 PO-Flow for Potential Outcomes Prediction
@@
Potential Outcomes Prediction
Algorithm 1
Potential Outcomes Prediction
To sample potential outcomes $\hat{y}^{(a)}$ for an individual $(x,a)$, the procedure follows a simple reverse process, as defined in Equation (11). Starting from a noise sample $z\sim q(\cdot)$, we integrate the velocity field $v_{\theta}$ backward in time, conditioned on covariates $x$ and treatment $a$, to obtain a point estimate of the potential outcome $\hat{y}^{(a)}$. The steps are summarized in Algorithm 1.
In addition to predicting point estimates of potential outcomes, PO-Flow is capable of learning the densities of the potential outcomes, as detailed in Section 4.3.
4.2 PO-Flow for Counterfactual Outcomes Prediction
The counterfactual outcome prediction consists of both a forward and a reverse process. Given a factual $(y^{(a)},x,a)\sim p_{Y,X,A}(\cdot)$, the forward process (Lines 1–2 in Alg. 2) encodes $(y^{(a)},x,a)$ into a latent representation $z$. Subsequently, the reverse process (Lines 3–5) decodes $z$ under the counterfactual treatment 1a1-a, yielding the counterfactual prediction $\hat{y}_{\text{cf}}^{(1-a)}$. The process of counterfactual prediction is also illustrated in Figure 1.

### v3

# Flow-based Generative Modeling of Potential Outcomes and Counterfactuals

@@ Abstract
Predicting potential and counterfactual outcomes from observational data is central to individualized decision-making, particularly in clinical settings where treatment choices must be tailored to each patient rather than guided solely by population averages. We propose PO-Flow, a continuous normalizing flow (CNF) framework for causal inference that jointly models potential outcome distributions and factual-conditioned counterfactual outcomes. Trained via flow matching, PO-Flow provides a unified approach to individualized potential outcome prediction, conditional average treatment effect estimation, and counterfactual prediction. By encoding an observed factual outcome into a shared latent representation and decoding it under an alternative treatment, PO-Flow relates factual and counterfactual realizations at the individual level, rather than generating counterfactuals independently from marginal conditional distributions. In addition, PO-Flow supports likelihood-based evaluation of potential outcomes, enabling uncertainty-aware assessment of predictions. A supporting recovery guarantee is established under certain assumptions, and empirical results on benchmark datasets demonstrate strong performance across a range of causal inference tasks within the potential outcomes framework.
I Introduction
Predicting potential outcomes (POs) is central to causal inference and individualized decision-making. In the potential outcomes framework, each individual with covariates $X$ has two outcomes: $Y^{(1)}$ if treated ($A=1$) and $Y^{(0)}$ if untreated ($A=0)$, but only one factual outcome $Y^{(A)}$ is observed [12]. In applications such as healthcare, practitioners often require patient-level comparisons across treatments, rather than relying solely on population-level averages [9].
Most existing methods focus on estimating the conditional average treatment effect (CATE), $\mathbb{E}[Y^{(1)}-Y^{(0)}\mid X]$, using meta-learners and model-specific approaches such as representation learning and deep models [18, 7, 8, 2, 29]. However, strong CATE performance does not necessarily imply accurate individualized potential outcome prediction in finite samples, since the assumptions are sufficient for CATE estimation can be weaker than those required for individual-level modeling [28].
@@ I Introduction
Accurately predicting individualized potential outcomes and learning their conditional distributions $p(Y^{(a)}\mid X)$ for each treatment $A=a$ is therefore a crucial task, particularly for uncertainty quantification and risk-sensitive decision-making [15]. Recent work has explored generative models for sampling potential outcomes and estimating empirical distributions [38, 23, 24]. Nevertheless, methods that learn the full density of potential outcomes within the potential outcomes framework, without relying on restrictive distributional assumptions (e.g., Gaussian mixtures), are still relatively limited in the literature.
Beyond potential outcome prediction, an important but less-studied causal query is factual-conditioned counterfactual prediction,
$Y_{\mathrm{cf}}^{(1-A)}:=Y^{(1-A)}\mid X,\,Y^{(A)},$
which asks what would have happened under the alternative treatment, given the observed factual outcome. This differs from standard PO prediction $p(Y^{(A)}\mid X)$, which does not condition on the realized outcome. Conditioning on $Y^{(A)}$ induces dependence between factual and counterfactual outcomes and is especially relevant for post-hoc analyses (e.g., “would this patient have done better on another therapy?”) [37, 10]. Structural causal models (SCMs) provide a canonical framework for such reasoning, but typically require specifying a causal graph and parametric structural equations [26, 5, 27].
In contrast, within the potential outcomes literature, general-purpose methods for individualized factual-conditioned counterfactual prediction remain underexplored. [^1] Notably, some works [16, 21, 19, 38] use the term “counterfactual” to refer to the unobserved potential outcome under an alternative treatment, without conditioning on the observed factual outcome. This motivates the need for methods that support counterfactual reasoning without requiring a parametric SCM, while remaining compatible with the potential outcomes framework.
@@
Despite recent progress, two key gaps remain. First, counterfactual outcome estimation conditioned on the observed factual outcome has received limited attention, despite its importance for individualized and post-hoc analyses. Second, existing methods are often specialized—targeting CATE estimation, potential outcome prediction, or distributional modeling in isolation—rather than providing a unified and scalable framework that jointly addresses these tasks.
In this paper, we propose PO-Flow, a continuous normalizing flow (CNF) trained via flow matching [20], which provides a unified generative framework for
  • (i)
    individualized potential outcome prediction,
  • (ii)
    conditional average treatment effect (CATE) estimation,
  • (iii)
    factual-conditioned counterfactual prediction via an abduction–action–prediction procedure enabled by invertible flows, with a supporting recovery result (Corollary V.6), and
  • (iv)
    potential outcome density learning for likelihood-based evaluation and uncertainty quantification.
By encoding an observed factual outcome into a shared latent representation and decoding it under an alternative treatment, PO-Flow enables mapping an individual factual outcome to a corresponding counterfactual outcome at the realization level, rather than generating outcomes independently from marginal conditional distributions.
II Preliminaries
A Continuous Normalizing Flow (CNF) defines an invertible transformation between a data variable and a simple base variable via a time-dependent Neural Ordinary Differential Equation (ODE). Starting from the data $y(0)=y_{0}$ at time $t=0$, the trajectory $y(t)$ evolves according to
$\frac{dy(t)}{dt}=v_{\theta}(y(t),t),\quad t\in[0,1],$ (1)
@@ II Preliminaries
where $v_{\theta}:\mathbb{R}^{d}\times[0,1]\to\mathbb{R}^{d}$ denotes the velocity field parameterized by a neural network with parameters $\theta$.
Under standard regularity conditions, the Neural ODE induces an invertible mapping from the data $y_{0}$ at $t=0$ to a base variable $y_{1}=z$ at $t=1$. The inverse mapping from $y_{1}$ to $y_{0}$ is obtained by integrating (1) backward in time. Throughout this paper, we adopt the standard choice $y_{1}=z\sim\mathcal{N}(0,I)$ at $t=1$, and denote the corresponding base density by $q(\cdot)=N(0,I)$.
We train the CNF using conditional flow matching (CFM) [20, 4], which avoids explicit density simulation. Given a data sample $y_{0}$ and an independent base sample $y_{1}\sim q(\cdot)$, CFM defines a reference path $\phi_{t}=\phi(y_{0},y_{1};t)$ connecting $y_{0}$ and $y_{1}$. We adopt the linear interpolant
$\phi_{t}=(1-t)\,y_{0}+t\,y_{1},$ (2)
whose associated reference velocity is
$\frac{d\phi_{t}}{dt}=y_{1}-y_{0}.$ (3)
CFM trains the velocity field $v_{\theta}$ to match this reference velocity along the interpolation path by minimizing
@@
$\begin{split}&\mathcal{L}_{\mathrm{CFM}}(\theta)\\ =&~\mathbb{E}_{t\sim\mathcal{U}[0,1],\,y_{0}\sim p_{\text{data}},\,y_{1}\sim q(\cdot)}\left\|v_{\theta}(\phi_{t},t)-\frac{d\phi_{t}}{dt}\right\|^{2}.\end{split}$ (4)
After training, samples are generated by drawing $y_{1}\sim q(\cdot)$ and integrating (1) backward in time to obtain $y_{0}$.
III Problem Settings
In the Potential Outcomes (POs) framework, each individual is characterized by covariates $X\in\mathbb{R}^{d_{X}}$, a binary treatment assignment $A\in\{0,1\}$, and an observed factual outcome $Y^{(A)}\in\mathbb{R}$. We observe an i.i.d. dataset
$\mathcal{D}=\{(y_{i},x_{i},a_{i})\}_{i=1}^{n}\sim p_{Y,X,A},$
where $y_{i}=y_{i}^{(a_{i})}$ denotes the realized potential outcome under the received treatment.
@@ III Problem Settings
Each individual has two potential outcomes, $Y^{(0)}$ and $Y^{(1)}$, but only one is observed. Potential outcome (PO) prediction targets the conditional distribution $p(Y\mid X,A=a)$, which equals $p(Y^{(a)}\mid X)$ under Assumption III.1 when $A=a$. We use $Y^{(a)}$ and $\hat{Y}^{(a)}$ to denote the true and predicted potential outcomes under treatment $a$, respectively.
The counterfactual outcome asks what would have happened under the alternative treatment, given the observed factual outcome $Y^{(a)}$:
$Y_{\mathrm{cf}}^{(1-a)}:=Y^{(1-a)}\mid X,\,Y^{(a)}.$
We denote the true and predicted counterfactual outcomes by $Y_{\mathrm{cf}}^{(1-a)}$ and $\hat{Y}_{\mathrm{cf}}^{(1-a)}$, respectively.
@@
To ensure the identifiability of potential outcomes, we adopt the standard assumptions commonly used in the literature [22, 24, 38]:
Assumption III.1.
  1. 1.
    Consistency: If an individual receives treatment $A=a$, then $Y=Y^{(a)}$.
  2. 2.
    Unconfoundedness: All confounders are observed, i.e., $\{Y^{(0)},Y^{(1)}\}\perp\!\!\!\perp A\mid X$.
  3. 3.
    Overlap: $0<P(A=1\mid X=x)<1$ for all $x\in\mathcal{X}$.
@@
Later, we will introduce additional assumptions for factual-conditioned counterfactual prediction in PO-Flow and establish the corresponding theoretical properties.
@@
Algorithm 1 Potential Outcomes Prediction